# Imperfect quantum networks with tailored resource states

### v1

@@ III.3 One-dimensional cluster as a building block
When manipulating a cluster state into a Bell pair, after isolating the path, one is left with a 1D cluster with the target qubits at the ends, where the inner neighbors are measured in the corresponding basis to obtain a Bell pair. The number of qubits surrounding the path in higher-dimensional clusters depends on the shape of the chosen path, but as in Mor-Ruiz and Dür (2023), one can consider a straight-line path as a lower-bound approximation, as it has the maximum number of outer neighbors. This manipulation can be improved in terms of minimizing the number of required measurements as shown in Hahn et al. (2019) for the two-dimensional case, however, it can be used as a lower-bound scenario. Therefore, the 1D cluster is a well-suited building block for higher-dimensional clusters as resource states. Moreover, the use of path-finding algorithms is not necessary for the one-dimensional case. In the next results, we consider a 1D cluster as a resource state.
@@
Importantly, in Mor-Ruiz and Dür (2023) it is shown that cluster states of higher dimension under local depolarizing noise have better behavior as the size of the network increases. To analyze this noise on such clusters, in Mor-Ruiz and Dür (2023), the straight-line approximation is taken, such that that the impact of the depolarizing noise on the target Bell pair increases linearly with dimension, keeping the same structure as for the one-dimensional case, but exponentially with the number of inner neighbors. Nevertheless, in this work, we are also considering time-dependent dephasing noise. The measurement of the outer neighbors is insensitive to the dephasing noise, as they are always measured in the $\sigma_{z}$ basis. Thus, to understand the impact of the time-dependent dephasing noise, on a high-dimensional cluster is sufficient to study the impact on the inner neighbors and target qubits.
@@
For a specific geometry, the distribution of the 1D cluster is not uniquely defined, such that it can be customized. For example, considering the process of generation of the resource state, we assume that the quantum communication channels are established between neighboring stations, so the most basic choice is to place the neighboring qubits in the 1D cluster in neighboring network nodes, then, entanglement does not have to be distributed over long distances. We refer to this distribution of the 1D cluster as the basic entanglement topology, depicted in Fig. 3a.
@@
However, one can customize the entanglement topology with the targets of the network in mind. If the pairs of nodes that are most probable to request a Bell state are known, one can distribute the resource state such that these pairs have a direct entanglement link or a minimal number of qubits in between in the entanglement topology. Then, the number of measurements in the manipulation to achieve these pairs is minimized. An instance of this entanglement topology is depicted in Fig. 3b. Later in this work, we analyze the advantages of having a target-oriented custom entanglement topology in terms of the fidelities of the target Bell pairs. Despite the possible advantages of minimizing the noise impact, such customization would require more advanced techniques for the entanglement generation process.
(a) Basic
(b) Custom
@@
Fig. 3: Entanglement topologies of a 5-qubit 1D cluster. The squares are the nodes, the snake-like lines are the quantum communication channels, the circles are qubits and the solid straight lines represent the entanglement between qubits. The network nodes and the quantum communication channels in (a) and (b) are the same.
IV Protocols
@@
Since, in this work, we discuss distributed graph states with qubits stored in imperfect quantum memories at spatially separated parties, the use of the one-shot approach is crucial. With this approach, measurements are performed as soon as possible, as there is no need to wait for any measurement outcome. Therefore, this avoids potential additional delays in network protocols, particularly if the geometry and entanglement topology do not match. We describe two protocols based on different structures and capabilities of the network nodes.
IV.1 Local protocol
The local protocol we consider is characterized by decision-making nodes. We assume that all nodes have the same capacities and they organize the manipulation between themselves. The structure of the classical communication channels here follows a one-dimensional chain. So, each node has two classical communication channels that connect it to the two closest nodes, but the ends of the chain only have one, as depicted in Fig. 4a. Such that the communication between nodes is done through this chain. Therefore, if the geometry and topology do not coincide, for a message to arrive at a certain node, it may have to pass through other nodes that might not be involved in the manipulation.
For the nodes to implement this coordination, after the dynamic phase of the network, they agree on a measurement pattern to achieve a network request, a so-called rule set using classical communication. Furthermore, all nodes know the geometry and the entanglement topology of the network, so each of them knows if it is involved in the manipulation, and what action it should take.
Local Protocol
Input: One of the target nodes, assume $a$, requires a Bell pair with $b$. 1. Node $a$ forwards this request to the closest nodes and initializes the chain of measurements: (a) Once the closest nodes receive the request, they decide if the message should be relayed to their other closest nodes or not. This decision depends on whether they are on the path of classical communication to one of the qubits in $\{q_{i}\}$, which are defined by the entanglement topology. (b) These nodes check the entanglement topology to see if their qubit belongs to $\{q_{i}\}$, and if so, consult the rule set to determine the basis of the measurement. (c) This relay-measure procedure continues until no further relying is necessary. Meanwhile, the nodes that have performed a measurement send the outcome to the corresponding target nodes via the classical channels. 2. The target nodes receive all outcomes. 3. $a$ and $b$ apply the corrections Output: $a$ and $b$ share a Bell pair.
@@ IV.1 Local protocol
(a) Local
(b) Central
Fig. 4: Location of the classical communication channels for the local and the central protocols for a 5-node network. The squares are the nodes, which are the same for (a) and (b), and the double lines are the classical communication channels. The coordinator in (b) is the bottom-right node.
IV.2 Central protocol
@@
A request enters the network through the coordinator who is in charge of organizing the manipulation to achieve the target. Therefore, the classical communication channels of this protocol are established from each node to the coordinator, following a star-like structure as depicted in Fig. 4b.
Central Protocol
Input: Arrival of the request at $C$. 1. $C$ chooses the measurement pattern to achieve the target pair. 2. $C$ prepares the commands for $\{n_{i}\}$, which include (a) basis of the measurement ($\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$), (b) order to perform the measurement, (c) order to send back the outcome. 3. $C$ sends the commands to $\{n_{i}\}$ classically, through the classical channel between $C$ and each party. 4. Given the one-shot approach, as each of the qubits in $\{n_{i}\}$ receives the command, they perform it. 5. When $C$ has received all the outcomes, the correction operators are computed and sent to $a$ and $b$. 6. $a$ and $b$ apply the corrections Output: $a$ and $b$ share a Bell pair.
@@
V Simulation results
Here we present the results obtained from the simulation and optimization of specific scenarios of the setting presented in Sec. III with the protocols described in Sec. IV. For all results, we consider that classical information travels at the speed of light in optical fiber, $2\cdot 10^{8}$ m/s, and that all nodes have a processing time of 1 $\upmu$s to perform measurements on their qubit. The parameters that change in each studied scenario are the node positioning, the strength of the depolarizing noise, $p$, and the dephasing times, $T_{v}$. In Appendix A.4, we detail how the noise of the measurement patterns is studied in the simulation.
V.1 Optimal measurement pattern
@@
We label the qubits and nodes as $\{1,\dots,N\}$, where the targets are 1 and $N$ and the inner neighbors are $\{2,\dots,N-1\}$. Importantly, in the local protocol, we assume that node 1 is the one starting the protocol, and in the central protocol, we assume that the coordinator is node $\left\lceil N/2\right\rceil$. We express a measurement pattern as $\boldsymbol{m}$, an $N-2$ length vector where $m_{i}\in\{\sigma_{x},\sigma_{y},\sigma_{z}\}$, and the $i^{\text{th}}$ vector element corresponds to the measurement basis on the $(i-1)^{\text{th}}$ qubit.
The results show a binary behavior of the optimal measurement pattern, such that we can distinguish two regimes. One where the depolarizing noise is dominant and another where the dephasing noise is dominant, which corresponds to low $T$. Each regime has an optimal measurement pattern that changes when entering the other regime. Importantly, there is an area in between where the two measurement patterns are optimal and equivalent, which we call the transitional regime.
If $N$ is even, in the depolarizing regime the optimal measurement pattern is $(\sigma_{x},\sigma_{x},\dots,\sigma_{x})$, for both protocols. However, for the dephasing regime, the optimal measurement pattern is different. For the local protocol is $(\sigma_{y},\sigma_{y},\sigma_{x},\dots,\sigma_{x})$, and for the central protocol the optimal measurement pattern has the following elements:
@@
(a) Local protocol
(b) Central protocol
Fig. 5: Difference between the fidelities achieved using the optimal measurement pattern in the depolarizing regime ($F_{\mathcal{D}}$), and the dephasing regime ($F_{\mathcal{Z}}$), for both local and central protocols, with even $N$, in terms of the strength of the depolarizing noise $p$ and the dephasing time $T$. The striped area denotes that both fidelities $F_{\mathcal{D}}$ and $F_{\mathcal{Z}}$ are below 0.5, and the target Bell pair is no longer entangled.
@@ V.2 Introduction of asymmetry
V.2.1 Asymmetry in memory quality
@@
(a) Local protocol
(b) Central protocol
Fig. 6: Fidelity of a Bell pair obtained from a 7-qubit 1D cluster in terms of the dephasing time of the faulty memory with the local or the central protocol. Each line denotes which is the faulty qubit. Lines that coincide are plotted in the same color.
In Fig. 6, we plot the fidelity of the final Bell pair in terms of the dephasing time of the faulty memory, for both protocols with $N=7$, where we observe the behavior described above. For the local protocol, a faulty memory in one of the target qubits (qubits 1 and 7 in Fig. 6a) has the same effect, as they have to wait for each other, furthermore, since their delay times are longer than the others the impact of their faulty memory is higher. For the central protocol, we see that if the coordinator (qubit 4 in Fig. 6b) has a faulty memory it does not affect the fidelity, as the qubit in the coordinator is only subject to the delay time related to the processing time. Also, qubits that have the same distance to the coordinator have equal delay times, such that the change in fidelity due to a faulty memory in these qubits is the same, as one can see in the overlap of lines in Fig. 6b.
V.2.2 Asymmetry in position
@@
Fig. 7: Shifting of the position of node 4 in a 5-node network. The squares are the nodes, the circles are qubits and the lines represent the entanglement between them. Filled circles denote target qubits and empty circles denote inner neighbors.
@@
(a) Local protocol
(b) Central protocol
Fig. 8: Fidelity of a Bell pair obtained from a 5-qubit 1D cluster in terms of the shifted position of one of the nodes with the local or the central protocol. Each of the lines denotes which is the shifted qubit. The bottom $x$ axis denotes the distance to the previous node and the top is the corresponding distance to the next node. Lines that coincide in (a) are plotted in the same style and color.
@@
V.3 Custom resource state
@@
Fig. 9: Random positioning of 100 nodes on a flat surface, the line connecting dots denotes neighboring stations.
@@
(a) Local protocol
(b) Central protocol
Fig. 10: Average fidelity over the hops defined by the custom entanglement topology of a 100-qubit 1D cluster, for both the local and central protocols. The shadowed area shows the corresponding standard deviation.
@@
Fig. 11: Percentage of final pairs with fidelity greater or equal to 0.5 in terms of the hops defined by the custom entanglement topology of a 100-qubit 1D cluster, for both the local and central protocols.
@@ VI Conclusions and outlook
We have focused on simulating, optimizing, and tailoring the adaptive phase in imperfect EBQNs, where we consider a cluster state as a resource state and a Bell pair as a target state. In particular, we show that optimizing the measurement pattern used for this manipulation is not only relevant but also the optimal measurement pattern changes depending on the noise parameters in a symmetric scenario. We take this optimization further by considering asymmetries in the dephasing times of the qubits and the positioning of the network nodes, to find that the optimal measurement pattern does not change for the particular parameter space we analyze. Moreover, when considering these asymmetric scenarios we can quantify the impact of each asymmetry in terms of the fidelity of the target state. Lastly, we demonstrate that our method based on the NSF can consider large graph states (up to 100 qubits in our results, but also states of several thousand or tens of thousands of qubits can be simulated) in a highly asymmetric scenario. Here we show that tailoring both the measurement pattern and the entanglement topology to the type of requests that are expected in the network is of importance due to its advantage in terms of the noise on the final target state. In particular, the use of such tailored resource states has been shown to have a significant effect on (average) reachable fidelities.
@@
We have argued that the 1D cluster as a resource state is a building block for higher-dimensional clusters. However, to fully unleash the possibilities of EBQNs, considering higher-dimensional clusters is crucial, as the large connectivity of these clusters allows for parallel target states (this is limited in the one-dimensional instance). Nonetheless, establishing parallel connections in EBQNs is at its early stages, and only results in the noiseless case have been studied Freund et al. (2024), which makes it even more appealing to study in an imperfect setting as the one we use here. Nevertheless, in order to consider time-dependent noise with these states, the proposed protocols must be adapted to include network routing Caleffi (2017); Hahn et al. (2019); Leung et al. (2010); Gyongyosi and Imre (2017) (which is not needed for the one-dimensional instance), since choosing the shortest path in high-dimensional clusters in an asymmetric setting is not trivial. Moreover, network routing can also be required to expand the local protocol, e.g., considering that each node only knows about the nodes classically connected to it.
@@
Besides considering high-dimensional clusters, one can also investigate other families of graph states, e.g., tree states, which have been shown to have extremely good behavior under depolarizing noise Mor-Ruiz and Dür (2023) at the cost of destroying large parts of the cluster. Moreover, when considering other entanglement topologies one might also want to change the geometry of the network and explore other possibilities of the network protocols. Note that our approach can be used as an inspiration for these new protocols and, moreover, our simulation tools can be freely used to analyze them.
@@
The code we have used for all the results in this work is archived at Mor-Ruiz and Wallnöfer (2024). An implementation of the NSF for graph states is available as a Python package at Wallnöfer (2024).

### v2

@@ 3.3 One-dimensional cluster as a building block
When manipulating a cluster state into a Bell pair, after isolating the path, one is left with a 1D cluster with the target qubits at the ends, where the inner neighbors are measured in the corresponding basis to obtain a Bell pair. The number of qubits surrounding the path in higher-dimensional clusters depends on the shape of the chosen path, but as in [21], one can consider a straight-line path as a lower-bound approximation, as it has the maximum number of outer neighbors. This manipulation can be improved in terms of minimizing the number of required measurements as shown in [28] for the two-dimensional case, however, it can be used as a lower-bound scenario. Therefore, the 1D cluster is a well-suited building block for higher-dimensional clusters as resource states. Moreover, the use of path-finding algorithms is not necessary for the one-dimensional case. In the next results, we consider a 1D cluster as a resource state.
@@
Importantly, in [21] it is shown that cluster states of higher dimension under local depolarizing noise have better behavior as the size of the network increases. To analyze this noise on such clusters, in [21], the straight-line approximation is taken, such that the impact of the depolarizing noise on the target Bell pair increases linearly with dimension, keeping the same structure as for the one-dimensional case, but exponentially with the number of inner neighbors. Nevertheless, in this work, we are also considering time-dependent dephasing noise. The measurement of the outer neighbors is insensitive to the dephasing noise, as they are always measured in the $\sigma_{z}$ basis. Thus, to understand the impact of the time-dependent dephasing noise, on a high-dimensional cluster is sufficient to study the impact on the inner neighbors and target qubits.
@@
For a specific geometry, the distribution of the 1D cluster is not uniquely defined, such that it can be customized. For example, considering the process of generation of the resource state, we assume that the quantum communication channels are established between neighboring stations, so the most basic choice is to place the neighboring qubits in the 1D cluster in neighboring network nodes, then, entanglement does not have to be distributed over long distances. We refer to this distribution of the 1D cluster as the basic entanglement topology, depicted in Fig. 3(a).
@@
However, one can customize the entanglement topology with the targets of the network in mind. If the pairs of nodes that are most probable to request a Bell state are known, one can distribute the resource state such that these pairs have a direct entanglement link or a minimal number of qubits in between in the entanglement topology. Then, the number of measurements in the manipulation to achieve these pairs is minimized. An instance of this entanglement topology is depicted in Fig. 3(b). Later in this work, we analyze the advantages of having a target-oriented custom entanglement topology in terms of the fidelities of the target Bell pairs. Despite the possible advantages of minimizing the noise impact, such customization would require more advanced techniques for the entanglement generation process.
[…]
[width=0.5]Fig_3a.pdf
[…]
[width=0.5]Fig_3b.pdf
@@
Figure 3: Entanglement topologies of a 5-qubit 1D cluster. The squares are the nodes, the snake-like lines are the quantum communication channels, the circles are qubits and the solid straight lines represent the entanglement between qubits. The network nodes and the quantum communication channels in (a) and (b) are the same.
4 Protocols
@@
Since, in this work, we discuss distributed graph states with qubits stored in imperfect quantum memories at spatially separated parties, the use of the one-shot approach is crucial. With this approach, measurements are performed as soon as possible, as there is no need to wait for any measurement outcome. Therefore, this avoids potential additional delays in network protocols, particularly if the geometry and entanglement topology do not match. We describe two protocols based on different structures and capabilities of the network nodes. Importantly, note that the following protocols only consider a single network request per resource state.
4.1 Local protocol
The local protocol we consider is characterized by decision-making nodes. We assume that all nodes have the same capacities and they organize the manipulation between themselves. The structure of the classical communication channels here follows a one-dimensional chain. So, each node has two classical communication channels that connect it to the two closest nodes, but the ends of the chain only have one, as depicted in Fig. 4(a). Such that the communication between nodes is done through this chain. Therefore, if the geometry and topology do not coincide, for a message to arrive at a certain node, it may have to pass through other nodes that might not be involved in the manipulation.
For the nodes to implement this coordination, after the dynamic phase of the network, they agree on a measurement pattern to achieve a network request, a so-called rule set using classical communication. Furthermore, all nodes know the geometry and the entanglement topology of the network, so each of them knows if it is involved in the manipulation, and what action it should take. Note that this rule set architecture is rather common in classical communication and has also been explored in the quantum regime, e.g., [66].
[…]
! Local Protocol Input: One of the target nodes, assume $a$, requires a Bell pair with $b$. 1. Node $a$ forwards this request to the closest nodes and initializes the chain of measurements: (a) Once the closest nodes receive the request, they decide if the message should be relayed to their other closest nodes or not. This decision depends on whether they are on the path of classical communication to one of the qubits in $\{q_{i}\}$, which are defined by the entanglement topology. (b) These nodes check the entanglement topology to see if their qubit belongs to $\{q_{i}\}$, and if so, consult the rule set to determine the basis of the measurement. (c) This relay-measure procedure continues until no further relying is necessary. Meanwhile, the nodes that have performed a measurement send the outcome to the corresponding target nodes via the classical channels. 2. The target nodes receive all outcomes. 3. $a$ and $b$ apply the corrections Output: $a$ and $b$ share a Bell pair.
@@ 4.1 Local protocol
[…]
[width=0.5]Fig_4a.pdf
[…]
[width=0.5]Fig_4b.pdf
Figure 4: Location of the classical communication channels for the local and the central protocols for a 5-node network. The squares are the nodes, which are the same for (a) and (b), and the double lines are the classical communication channels. The coordinator in (b) is the bottom-right node.
4.2 Central protocol
@@
A request enters the network through the coordinator who is in charge of organizing the manipulation to achieve the target. Therefore, the classical communication channels of this protocol are established from each node to the coordinator, following a star-like structure as depicted in Fig. 4(b).
[…]
! Central Protocol Input: Arrival of the request at $C$. 1. $C$ chooses the measurement pattern to achieve the target pair. 2. $C$ prepares the commands for $\{n_{i}\}$, which include (a) basis of the measurement ($\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$), (b) order to perform the measurement, (c) order to send back the outcome. 3. $C$ sends the commands to $\{n_{i}\}$ classically, through the classical channel between $C$ and each party. 4. Given the one-shot approach, as each of the qubits in $\{n_{i}\}$ receives the command, they perform it. 5. When $C$ has received all the outcomes, the correction operators are computed and sent to $a$ and $b$. 6. $a$ and $b$ apply the corrections Output: $a$ and $b$ share a Bell pair.
@@
5 Simulation results
Here we present the results obtained from the simulation and optimization of specific scenarios of the setting presented in Sec. 3 with the protocols described in Sec. 4. For all results, we consider that classical information travels at the speed of light in optical fiber, $2\cdot 10^{8}$ m/s, and that all nodes have a processing time of 1 $\upmu$s to perform measurements on their qubit. The parameters that change in each studied scenario are the node positioning, the strength of the depolarizing noise, $p$, and the dephasing times, $T_{v}$. In Appendix A.4, we detail how the noise of the measurement patterns is studied in the simulation.
5.1 Optimal measurement pattern
@@
We label the qubits and nodes as $\{1,\dots,N\}$, where the targets are 1 and $N$ and the inner neighbors are $\{2,\dots,N-1\}$. Importantly, in the local protocol, we assume that node 1 is the one starting the protocol, and in the central protocol, we assume that the coordinator is node $\left\lceil N/2\right\rceil$. We express a measurement pattern as $\bm{m}$, an $N-2$ length vector where $m_{i}\in\{\sigma_{x},\sigma_{y},\sigma_{z}\}$, and the $i^{\text{th}}$ vector element corresponds to the measurement basis on the $(i+1)^{\text{th}}$ qubit.
The results obtained for small systems (up to $N\leq 12$) show a binary behavior of the optimal measurement pattern, such that we can distinguish two regimes. One where the depolarizing noise is dominant and another where the dephasing noise is dominant, which corresponds to low $T$. Each regime has an optimal measurement pattern that changes when entering the other regime. Importantly, there is an area in between where the two measurement patterns are optimal and equivalent, which we call the transitional regime.
From the observed optimal patterns for small systems, we extrapolate the following measurement patterns for all $N$. If $N$ is even, in the depolarizing regime the optimal measurement pattern is $(\sigma_{x},\sigma_{x},\dots,\sigma_{x})$, for both protocols. However, for the dephasing regime, the optimal measurement pattern is different. For the local protocol is $(\sigma_{y},\sigma_{y},\sigma_{x},\dots,\sigma_{x})$, and for the central protocol the optimal measurement pattern has the following elements:
@@
[…]
[width=]Fig_5a.pdf
[…]
[width=]Fig_5b.pdf
Figure 5: Difference between the fidelities achieved using the optimal measurement pattern in the depolarizing regime ($F_{\mathcal{D}}$), and the dephasing regime ($F_{\mathcal{Z}}$), for both local and central protocols, with even $N$, in terms of the strength of the depolarizing noise $p$ and the dephasing time $T$. The striped area denotes that both fidelities $F_{\mathcal{D}}$ and $F_{\mathcal{Z}}$ are below 0.5, and the target Bell pair is no longer entangled.
@@ 5.2 Introduction of asymmetry
5.2.1 Asymmetry in memory quality
@@
[…]
[width=0.5]Fig_6a.pdf
[…]
[width=0.5]Fig_6b.pdf
Figure 6: Fidelity of a Bell pair obtained from a 7-qubit 1D cluster in terms of the dephasing time of the faulty memory with the local or the central protocol. Each line denotes which is the faulty qubit. Lines that coincide are plotted in the same color.
In Fig. 6, we plot the fidelity of the final Bell pair in terms of the dephasing time of the faulty memory, for both protocols with $N=7$, where we observe the behavior described above. For the local protocol, a faulty memory in one of the target qubits (qubits 1 and 7 in Fig. 6(a)) has the same effect, as they have to wait for each other, furthermore, since their delay times are longer than the others the impact of their faulty memory is higher. For the central protocol, we see that if the coordinator (qubit 4 in Fig. 6(b)) has a faulty memory it does not affect the fidelity, as the qubit in the coordinator is only subject to the delay time related to the processing time. Also, qubits that have the same distance to the coordinator have equal delay times, such that the change in fidelity due to a faulty memory in these qubits is the same, as one can see in the overlap of lines in Fig. 6(b).
5.2.2 Asymmetry in position
@@
[…]
[width=]Fig_7.pdf
Figure 7: Shifting of the position of node 4 in a 5-node network. The squares are the nodes, the circles are qubits and the lines represent the entanglement between them. Filled circles denote target qubits and empty circles denote inner neighbors.
@@
[…]
[width=0.5]Fig_8a.pdf
[…]
[width=0.5]Fig_8b.pdf
Figure 8: Fidelity of a Bell pair obtained from a 5-qubit 1D cluster in terms of the shifted position of one of the nodes with the local or the central protocol. Each of the lines denotes which is the shifted qubit. The bottom $x$ axis denotes the distance to the previous node and the top is the corresponding distance to the next node. Lines that coincide in (a) are plotted in the same style and color.
@@
5.3 Custom resource state
@@
[…]
[width=]Fig_9.pdf
Figure 9: Random positioning of 100 nodes on a flat surface, the line connecting dots denotes neighboring stations.
@@
[…]
[width=]Fig_10a.pdf
[…]
[width=]Fig_10b.pdf
Figure 10: Average fidelity over the hops defined by the custom entanglement topology of a 100-qubit 1D cluster, for both the local and central protocols. The shadowed area shows the corresponding standard deviation.
@@
[…]
[width=]Fig_11.pdf
Figure 11: Percentage of final pairs with fidelity greater or equal to 0.5 in terms of the hops defined by the custom entanglement topology of a 100-qubit 1D cluster, for both the local and central protocols.
@@ 6 Conclusions and outlook
We have focused on simulating, optimizing, and tailoring the adaptive phase in imperfect EBQNs, where we consider a cluster state as a resource state and a Bell pair as a target state. In particular, we show that optimizing the measurement pattern used for this manipulation is not only relevant but also the optimal measurement pattern changes depending on the noise parameters in a symmetric scenario. We take this optimization further by considering asymmetries in the dephasing times of the qubits and the positioning of the network nodes, to find that the optimal measurement pattern does not change for the particular parameter space we analyze. Moreover, when considering these asymmetric scenarios we can quantify the impact of each asymmetry in terms of the fidelity of the target state. Lastly, we demonstrate that our method based on the NSF can consider large graph states (up to 100 qubits in our results, but also states of several thousand or tens of thousands of qubits can be simulated) in a highly asymmetric scenario, given our noise model and resource state. Here we show that tailoring both the measurement pattern and the entanglement topology to the type of requests that are expected in the network is of importance due to its advantage in terms of the noise on the final target state. In particular, the use of such tailored resource states has been shown to have a significant effect on (average) reachable fidelities.
@@
We have argued that the 1D cluster as a resource state is a building block for higher-dimensional clusters. However, to fully unleash the possibilities of EBQNs, considering higher-dimensional clusters is crucial, as the large connectivity of these clusters allows for parallel target states (this is limited in the one-dimensional instance). Nonetheless, establishing parallel connections in EBQNs is at its early stages, and only results in the noiseless case have been studied [67], which makes it even more appealing to study in an imperfect setting as the one we use here. Nevertheless, in order to consider time-dependent noise with these states, the proposed protocols must be adapted to include network routing [68, 28, 69, 70] (which is not needed for the one-dimensional instance), since choosing the shortest path in high-dimensional clusters in an asymmetric setting is not trivial. Moreover, network routing can also be required to expand the local protocol, e.g., considering that each node only knows about the nodes classically connected to it. Additional networking aspects, such as which requests of target states can be fulfilled simultaneously, would need to be considered in this case.
@@
Besides considering high-dimensional clusters, one can also investigate other families of graph states, e.g., tree states, which have been shown to have extremely good behavior under depolarizing noise [21] at the cost of destroying large parts of the cluster. Moreover, when considering other entanglement topologies one might also want to change the geometry of the network and explore other possibilities of the network protocols. Note that our approach can be used as an inspiration for these new protocols and, moreover, our simulation tools can be freely used to analyze them.
@@
The code we have used for all the results in this work is archived at [71]. An implementation of the NSF for graph states is available as a Python package at [57].